OpenStudy (anonymous):

help \[\Huge \frac{ 1 }{ 1+\frac{ 1 }{ 1-\frac{ 1 }{ 1+\frac{ 1 }{ x } } } }\]

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+\frac{ 1 }{ 1-\frac{ 1 }{ 1+\frac{ 1 }{ x } } } }\]

OpenStudy (anonymous):

help people

OpenStudy (anonymous):

complex fraction here

ganeshie8 (ganeshie8):

\[\Huge \frac{ 1 }{ 1+\frac{ 1 }{ 1-\frac{ 1 }{ \color{Red}{1+\frac{ 1 }{ x }} } } }\]

ganeshie8 (ganeshie8):

get common denominator and add those two first

ganeshie8 (ganeshie8):

\[\Huge \frac{ 1 }{ 1+\frac{ 1 }{ 1-\frac{ 1 }{ \color{Red}{\frac{x}x{}+\frac{ 1 }{ x }} } } }\]

ganeshie8 (ganeshie8):

\[\Huge \frac{ 1 }{ 1+\frac{ 1 }{ 1-\frac{ 1 }{ \color{Red}{\frac{x+1}x} } } }\]

ganeshie8 (ganeshie8):

keep simplifying...

OpenStudy (anonymous):

???????????????

OpenStudy (dan815):

yep carry on his work do the new common deominator and continue

OpenStudy (anonymous):

how is that?

OpenStudy (dan815):

okok

OpenStudy (dan815):

\[\large {1 - \frac{1}{\frac{x+1}{x}}}\]

OpenStudy (dan815):

simplify that part

OpenStudy (anonymous):

2x + 1

OpenStudy (anonymous):

?